OpenStudy (anonymous):

Can someone help me please i really need the answer -F(x)=x+6 G(x)=12x-7 what is F(x) + G(x)? A. 12x-1 B. 13x-1 C. 13x+13 D. 12x + 13

OpenStudy (ybarrap):

add everything to the right of the equal signs to together

OpenStudy (anonymous):

how would i do that though? x+6+12x-7? then what

OpenStudy (ybarrap):

This is good, now look at each term individually and group them according to if it has an "X" next to it or not

OpenStudy (anonymous):

how do i bring the x to the other side though?

OpenStudy (ybarrap):

in your equation x+6+12x-7, you have two terms with "x": x + 12x. Right?

OpenStudy (anonymous):

oh yes i do

OpenStudy (ybarrap):

you also have two terms that do not have "x": 6 -7. Now what does this equal?

OpenStudy (ybarrap):

x+12x = ? 6 - 7 = ?

OpenStudy (anonymous):

oh okay and it equals 1and x + 12x =13x?

OpenStudy (ybarrap):

close, 6-7 = -1 So now we have x+12x = 13x and 6-7 = -1 Next we combine. You had x+6+12x-7 Then we found x+12x + 6 -7 = 13x - 1 That's it!

OpenStudy (anonymous):

oh wow thanks for your help i appreciate it since i have a test tommarow morning! would you be able to help me with any other ones?

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

yes